### v2

# Physics-Inspired Synthesized Underwater Image Dataset

Abstract
This paper introduces the physics-inspired synthesized underwater image dataset (PHISWID), a dataset tailored for enhancing underwater image processing through physics-inspired image synthesis. For underwater image enhancement, data-driven approaches (e.g., deep neural networks) typically demand extensive datasets, yet acquiring paired clean and degraded underwater images poses significant challenges. Existing datasets have limited contributions to image enhancement due to lack of physics models, publicity, and ground-truth images. PHISWID addresses these issues by offering a set of paired ground-truth (atmospheric) and underwater images synthetically degraded by color degradation and marine snow artifacts. Generating underwater images from atmospheric RGB-D images based on physical models provides pairs of real-world ground-truth and degraded images. Our synthetic approach generates a large quantity of the pairs, enabling effective training of deep neural networks and objective image quality assessment. Through benchmark experiment with some datasets and image enhance methods, we validate that our dataset can improve the image enhancement performance. Our dataset, which is publicly available, contributes to the development in underwater image processing.
Index Terms:
underwater image enhancement, marine snow, creating dataset, deep learning.
I Introduction
Image restoration and enhancement [2, 3, 4, 5, 6, 7, 8] has been a long-studied topic for decades, with substantial advancements in the deep learning era. This progress is largely due to the availability of extensive datasets comprising pairs of ground-truth and degraded images. However, the situation changes dramatically when dealing with extreme image restoration scenarios, including underwater, satellite, and medical imaging [9, 10, 11, 1, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. These contexts present atypical degradation patterns that diverge from conventional Gaussian models, and they suffer from a scarcity of training images for algorithm refinement, particularly for deep neural networks.
Underwater image enhancement is one of the extreme image restoration challenges. The primary hurdle lies in the absence of ground-truth (atmospheric) images. Unlike image processing in atmospheric conditions, obtaining clean atmospheric counterparts for underwater images is not feasible. This absence of ground truth makes it difficult to objectively assess restoration quality. Furthermore, the lack of pairs of ground-truth and degraded images hampers neural network training, as traditional loss functions (e.g., peak signal to noise ratio (PSNR) and structual similarity (SSIM)) evaluate the error between these pairs [23].
Various datasets and enhancement methods have tackled to address this lack of ground-truth images[24, 1, 13, 25, 14, 26, 27, 28]. However, most datasets are not publicly available and are often oversimplified color degradation processes (details are described in Section II-A). Moreover, they overlook a critical aspect of degradation: Marine snow. Marine snow, visible as white dots in digital images as shown in Fig. 1, is caused by light reflecting off particles like organic matter and sand. This phenomenon degrades overall image quality, yet its removal techniques are still nascent, often limited to classical median-filtering approaches [29, 30, 31]. The lack of datasets has also made it difficult to develop effective removal techniques for marine snow.
In this paper, we propose a new dataset, the physics-inspired synthesized underwater image dataset (PHISWID), providing pairs of ground-truth images, and underwater images with color degradation and marine snow artifacts [32, 33]. Our dataset is based on physical observation models which enable to synthesis underwater image from real-world atmospheric images. For the physical observation models, the light absorption and scattering model in underwater is used to model color shift [34, 35] and the light scattering model for underwater particle to marine snow [36]. We apply these models to an atmospheric RGB-D image to generate its synthetic underwater image, obtaining a pair of the ground-truth and underwater images.
Our dataset is validated with various image enhancement methods and some existing datasets. We show that the learning-based methods trained with our dataset outperforms the ones trained with the other datasets [11, 37], even with basic U-Net [38]. Our benchmark results suggest a potential contribution of our dataset, which is made publicly available, to development of the underwater image processing [^1] https://github.com/reina0112/PHISMID-PHISWID..
Figure 1: Marine snow artifacts in real underwater images.
TABLE I: List of underwater image datasets.
Dataset Target Number of Images Method Availability
PHISWID (proposed) Color shift & Marine snow 2264 Physics-based synthesis
Dataset used in [25] Color shift & Haze 10143 Physics-based synthesis & Filtering
UIEB [11] Color shift 860 Voting
LSUI [37] Color shift 4279 Voting
SAUD [39] Color shift 100 Voting
EUVP [40] Color shift 12000 Cycle GAN
Dataset used in [12] Color shift 200 Physics-based synthesis
Dataset used in [41] Color shift 1449 Physics-based synthesis
MSRB [42] Marine snow 2700 Observation-based synthesis
The remainder of this paper is organized as follows: Section II reviews related studies for underwater image enhancement. We propose underwater image synthesis methods along with the marine snow observation model and the color shift model in Section III. The PHISWID specifications are introduced in Section IV. The benchmarking results with PHISWID are shown in Section V. Real-world examples and limitations are described in Section VI. Finally, we provide some concluding remarks in Section VII.
II Related Work
In this section, we introduce previous works on underwater image enhancement. We first introduce underwater image enhancement and available datasets, then focus on marine snow removal. Currently available datasets are listed in Table I.
II-A Underwater Image Enhancement and Available Datasets
The current trend of underwater image enhancement is based on deep learning as with the other image enhancement/restoration tasks[43, 11, 37, 44, 45]. In [43], a wavelength-attributed deep neural network is proposed. In [37], a transformer designed for underwater image enhancement is presented. A color correction module followed by an optical model is used in [11, 28]. These methods were trained with datasets including pairs of ground-truth (clean) and underwater (degraded) images for underwater image enhancement. However, collecting extensive pairs in underwater situations is still a challenge. This limits the performance of deep learning-based techniques.
An approach for creating the image pairs is that a clean (ground-truth-like) image is selected among images processed by various enhancement methods by human voting. We call it voting-based approach. UIEB dataset [11], which is widely utilized for developing many enhancement methods [43, 11], provides 860 real-world underwater images and corresponding ground-truth-like images across various water conditions. LSUI dataset [37] is also widely used [37, 46]. It contains 4279 real-world underwater images and the corresponding reference images. SAUD dataset [39] contains 100 raw underwater images and enhanced images used some existing methods. In the voting-based datasets, there is a possibility that no enhancement method works well on a certain underwater image. Therefore, they do not guarantee that the clean image is close to the true ground-truth image which is taken under the atmospheric environment. This gap in the ground-truth images could ruin the reliability of the datasets and lead to limited performance.
Generative adversarial network (GAN)-based approach [14, 47, 48, 40] is also used to generate clean, ground-truth-like images. The GANs are trained with sets of clean and degraded underwater images. A critical limitation of this approach, which does not provide pairs of ground-truth and degraded images, is that well-known full-reference image quality metrics like PSNR and SSIM cannot be computed to evaluate the enhancement/restoration accuracy. Moreover, many datasets are not publicly available and their contribution to image restoration development is limited. Furthermore, there is no dataset including both marine snow artifacts and color shift degradation.
Unlike the voting-based and GAN-based approaches, which generate the ground-truth-like images from underwater images, there is an approach to generate underwater images from ground-truth images which are taken under the atmospheric environment. We call it physics-based synthesis approach. In this approach, degraded underwater images are generated from atmospheric images by simulating the degradation processes of absorption and scattering [12, 41, 1, 25]. This approach requires precise models of degradation. However, some datasets [12, 41] are based on oversimplified color degradation processes. They ignore the light scattering and the radiance ratio of the scene reaching the camera. This synthesis may be incomplete for underwater image enhancement purposes. Furthermore, they are not available online (see Table I). It leads to a limited contribution.
II-B Marine Snow Removal
Marine snow is one of the main sources of underwater image degradation but has been underaddressed so far. Here, we briefly review some existing marine snow removal approaches.
A widely used method for marine snow removal is a median filter (MF) [49, 50, 29, 51]. If marine snow artifacts are assumed sufficiently small (typically 11113333 pixels in diameter), MF works well since small marine snow artifacts are similar to salt-and-pepper noise. However, as shown in Fig. 1, marine snow artifacts often become large. As we set a large filter size for MF to remove such artifacts, MF significantly blurs the entire image.
So far, a few deep learning-based marine snow removal methods have been proposed [26, 42, 52]. A typical training dataset consists of underwater images and their corresponding version with synthesized marine snow artifacts. In [26], marine snow particles are manually appended to underwater images with Adobe Photoshop. This results in a lack of randomness and scalability. In [42], shapes of synthesized marine snow artifacts in the dataset are modeled with ellipses. However, this marine snow synthesis is not based on a physical measurement model of a particle in underwater and the reproduction accuracy of marine snow is limited.
As a result, for underwater image enhancement, we need a dataset including the following images:
  • Ground truth clean atmospheric images, i.e., target images.
  • Underwater images corresponding to the atmospheric images degraded by absorption, scattering, and marine snow.
Creating such a dataset by taking many pictures in the real world in various water types is infeasible. Hence, we address this problem by synthesizing images based on physics-guided models for color shift and marine snow.
III Underwater Image Synthesis
Figure 2: Flow of underwater image synthesis for PHISWID. First, RGB and depth images are used to synthesize underwater images based on [1]. Then, marine snow artifacts are added with the method introduced in Section III.
In this section, we describe our method for generating underwater images from clean RGB-D atmospheric images by synthesizing degradation caused by absorption, scattering, and marine snow.
The flow for image synthesis is illustrated in Fig. 2. We first synthesize color-shifted images based on a physics-inspired color shift model proposed in Section III-A through the procedure described in Section III-B. Then, we also synthesize marine snow artifacts based on a model proposed in Section III-C through the procedure described in Section III-D.
III-A Color Shift Model
Here, we first describe the models simulating color shift of underwater images.
III-A1 Algorithm Overview
Algorithm 1 Color shift synthesis algorithm
0:  Clean RGB-D images : $D_{\text{RGB-D}}$ttenuation coefficients for each water type $w$ : $D_{w,\beta(\lambda)}$amera spectral response : $D_{\mathrm{SR}(\lambda),\text{camera}}$llumination irradiance on the water surface : $\Theta_{0}(\lambda)$$d_{\text{vert},\text{max}}$, $B_{c,\text{min}}$, $B_{c,\text{max}}$, $\rho(\lambda)=1$
0:  Synthesized image set corresponding : $D_{\mathcal{W},\text{synthesized}}$
1:  $\mathcal{W}$leftarrow {I, IA, IB, II, III, 1C, 3C, 5C, 7C, 9C}
2:  for $t=1$ to $10$ do
3:     $\beta(\lambda)$leftarrow $D_{\mathcal{W}_{t},\beta(\lambda)}$
4:     repeat
5:        $\Phi$leftarrow Clean image in $D_{\text{RGB-D}}$
6:        $z$leftarrow Depth map in $D_{\text{RGB-D}}$
7:        Adjust the dynamic range of $D_{h}$ into $d_{\text{horiz}}\in[0.5,14]$
8:        for $k=1$ to $10$ do
9:           Set $d_{\text{vert}}$ within $0.5\leq d_{\text{vert}}\leq d_{\text{vert},\text{max}}$ randomly
10:           Set $B_{c}$ within $B_{c,\text{min}}\leq B_{c}\leq B_{c,\text{max}}$ randomly
11:           Randomly select a camera $r$
12:           $\mathrm{SR}(\lambda)$leftarrow $D_{\mathrm{SR}(\lambda),r}$
13:           Calculate equation (1)–(5) for all pixels and get $U_{c}$
14:           Put $U_{c}$ into $D_{\mathcal{W}_{t},\text{synthesized}}$
15:        end for
16:     until Generating $U_{c}$ from all images in $D_{\text{RGB-D}}$
17:  end for
TABLE II: Ten Water types by Jerlov [34].
Ocean water type Coastal water type
I, IA, IB, II, III 1C, 3C, 5C, 7C, 9C
Low attenuation High attenuation
Our color shift model considers attenuation of light due to underwater condition and spectral difference of camera property. For modeling the attenuation, we use a model proposed by Jerlov [34] which presents attenuation of light at each wavelength (400 to 700 nm). The attenuation coefficients are varied for water types, which Jerlov classified into ten as shown in Table II. In addition to the attenuation property, the color also depends on the spectrum response which is an inherent property of a camera and lens. Therefore, our color shift synthesis takes into account the attenuation and spectral response to simulate the color shift in underwater environment.
Algorithm 1 describes the overview of our color shift synthesis method. The required inputs are listed as follows:
  • Clean RGB-D images: $D_{\text{RGB-D}}$.
  • Set of attenuation coefficients $\beta(\lambda)$ for each water type $w$, in which $\lambda$ is the wavelength: $D_{w,\beta(\lambda)}$.
  • Set of RGB spectral responses $\mathrm{SR}(\lambda)$ for several digital cameras: $D_{\mathrm{SR}(\lambda),\text{camera}}$.
  • Illumination irradiance on the water surface: $\Theta_{0}(\lambda)$.
  • Maximum vertical distance from water surface [^2] In this paper, we do not use “depth” for water depth not to confuse readers.: $d_{\text{vert},\text{max}}$
  • Maximum and minimum values of the background light: $B_{c,\text{max}}$ and $B_{c,\text{min}}$.
As a result, our algorithm creates ten sets of synthetic images, each corresponds to one water type.
III-A2 Underwater Image Degradation Model
In our color shift model, we consider two main sources of the degradation process: absorption and scattering. The underwater color shift model is represented as follows:
$U_{c}(x,y)=T_{c,d_{\text{vert}}}T_{c,d_{\text{horiz}}}\Phi(x,y)+B_{c}T_{c,d_{\text{vert}}}(1-T_{c,d_{\text{horiz}}}),$ (1)
where $c\in\{\text{R},\text{G},\text{B}\}$, $U_{c}(x,y)$ is the pixel values at $(x,y)$ of an underwater image, $\Phi(x,y)$ is the pixel values of the corresponding clean RGB-D image, $T_{c,d_{\text{vert}}}$ is the transmission map of color at the vertical distance $d_{\text{vert}}$, and $T_{c,d_{\text{horiz}}}$ is that in the horizontal direction (i.e., the distance between the camera and scene).
The first term in (1) represents the process of color absorption both in the vertical and the horizontal directions. The second term corresponds to the background color at the vertical distance $d_{\text{vert}}$ and the horizontal distance $d_{\text{horiz}}$. Further, $T_{c,d_{\text{vert}}}$ and $T_{c,d_{\text{horiz}}}$ is obtained by
$T_{c,d_{\text{vert}}}$ $=\text{exp}(-\beta_{c,d_{\text{vert}}}\cdot d_{\text{vert}})$ (2)
$T_{c,d_{\text{horiz}}}$ $=\text{exp}(-\beta_{c,d_{\text{horiz}}}\cdot d_{\text{horiz}}),$ (3)
where $\beta_{c,d_{\text{vert}}}$ and $\beta_{c,d_{\text{horiz}}}$ are the attenuation coefficients along the vertical and horizontal directions, respectively.
Based on [35], $\beta_{c,d}$ and $\beta_{c,d_{\text{horiz}}}$ is given as
[…]
where $\rho(\lambda)$ is the reflectance spectrum of the object surface. By using (4) and (5), we obtain attenuation coefficients to calculate $T_{c,d_{\text{vert}}}$ and $T_{c,d_{\text{horiz}}}$.
III-B Synthesizing Color Shift
By using the models introduced in Section III-A, we synthesize the color shift. We use the attenuation coefficients in all frequency ranges in contrast to the existing work [41] using only three coefficients.
The attenuation coefficients for ten water types $D_{w,\beta(\lambda)}$ (see Algorithm 1), are given by [53]. We also use the data of solar irradiance on the water surface [54] as $\Theta_{0}(\lambda)$. The camera spectrum responses are also given from the dataset of 28 digital cameras [55]. The other parameters are set to $d_{\text{vert},\text{max}}=1$, $\{B_{R,\text{min}},B_{G,\text{min}},B_{B,\text{min}}\}=\{0.4,0.7,0.7\}$, and $\{B_{R,\text{max}},B_{G,\text{max}},B_{B,\text{max}}\}=\{0.5,0.8,0.8\}$ [1]. We randomly determine five sets of $B_{c}$ and $d_{\text{vert}}$ from the uniform distribution in the range of $B_{c,\text{min}}\leq B_{c}\leq B_{c,\text{max}}$ and $0.5\leq d_{\text{vert}}\leq d_{\text{vert},\text{max}}$, respectively. As a result, 50 underwater images are generated from a single atmospheric image with the ten water types and the five sets of $B_{c}$ and $d_{\text{vert}}$.
III-C Marine Snow Model
III-C1 Light Scattering of Marine Snow
Although a precise model of marine snow artifacts requires to identify particles, the identification is impractical since any particle drifting in underwater could be the sources. Therefore, we consider light scattering in water for small particles based on the Jaffe-MacGlamery model [36, 56], which is common for underwater imagery [47, 57].
The Jaffe-MacGlamery model of a small object is illustrated in Fig. 3(a). The model considers direct-scattered and back-scattering components [36, 56]. Although the back-scattering components should be considered if a object is relatively large, we can ignore these components because objects caused marine snow should be small. Therefore, we simplify the model only with the direct-scattered components. Another simplification is that there is a single light source in a scene. We assume that a camera’s flash is the single source. The model with these two simplifications is illustrated in Fig. 3(b).
(a) Original Model
(b) Simplified Model
Figure 3: Jaffe-MacGlamery model.
TABLE III: Parameter settings for light scattering in (6) and (8).
Parameter Setting
$\theta$ 0
$E_{\mathrm{d}}(x,y)$ Direct-scattered component
$F_{l}$ Focal length of camera
$T_{l}$ Transmattance of the lens
$R$ Distance from camera to marine snow
$\beta_{c}$ Attenuation along the vertical and horizontal directions
$p(\cdot)$ Point spread function
The direct-scattered component $E_{\mathrm{d}}(x,y)$ at the coordinate $(x,y)$ in the observed image is thus represented as
$E_{\mathrm{d}}(x,y)=E_{\mathrm{I}}\left(x,y,0\right)e^{-\beta_{c}R}M\left(x,y\right)\frac{\cos^{4}\theta T_{l}}{4f_{n}}\left[\frac{R-F_{l}}{R}\right]^{2},$ (6)
where $E_{\mathrm{I}}\left(x,y,0\right)$ is defined as
$E_{\mathrm{I}}(x,y,0)=E^{\prime}_{\mathrm{I}}(x,y,0)*p(x,y|R,G,\beta_{c},B)+E^{\prime}_{\mathrm{I}}(x,y,0),$ (7)
and * is convolution, $M(x,y)$ is the reflectance values of the reflectance map for oceanographic objects. The other symbols are defined in Table III.
Here, we further assume the light source and the camera are located in the same 3D position, which results in $\theta=0$. This simplifies the observation model as in Fig. 3(b). By the above assumption, (6) can be rewritten as
$E_{\mathrm{d}}(x,y)=\left(A\frac{e^{-2\beta_{c}R}}{R^{2}}*p+A\frac{e^{-2\beta_{c}R}}{R^{2}}\right)\left[\frac{R-F_{l}}{R}\right]^{2},$ (8)
where $A$ is a constant coming from those in (6). The simplified model in (8) indicates that the light scattering of a marine snow particle only depends on $R$, i.e., the distance between the camera and the particle. The derivation of the equation is shown in Appendix A.
III-C2 Modeling Marine Snow Artifacts
Based on the observation model shown in Section III-C1, we model the type H and V333“H” and “V” refer to “highland” and “volcanic crater”, respectively, as in [42]. marine snow artifacts. The type H artifacts correspond to generic marine snow particles. The type V artifacts are often observed as an edge-enhanced version of the type H caused by automatic image processing in cameras (like those in smartphones). An example of type H and V marine snow is shown in Fig. 4.
We model a type H marine snow artifact as a Gaussian filtered version of a point source according to (8). The type V is the slightly modified version of the type H. Since the luminance of a particle depends on $R$, i.e., the distance between the camera and the particle, we model marine snow particles with the following steps:
  1. i.
    Determining coordinates of particles: As shown in Fig. 5, suppose that the camera is positioned parallel to the $(x,y)$-plane and at a sufficiently large distance from the scene located in the three-dimensional Euclidean space. The three-dimensional coordinates of the particles are uniform-randomly set. The number of marine snow particles is given for each image.
    A sparse matrix $\tilde{S}\in\mathbb{R}^{H\times W}$ for the particle coordinates, where $H$ and $W$ are the height and width of the target image, is created as
    $\tilde{S}(x,y)=\begin{cases}R&\text{if a particle is located at }(x,y)\\ 0&\text{otherwise,}\end{cases}$ (9)
    where $R$ is randomly set.
    We then make a set of matrices $\{E_{n}\}_{n}$ such that $\tilde{S}=\sum_{n=0}^{N-1}E_{n}$ based on $\tilde{S}$ as follows:
    $E_{n}(x,y)=\begin{cases}D&\text{if }d_{\text{horiz},n-1}<\tilde{S}(x,y)\leq d_{\text{horiz},n}\\ 0&\text{otherwise,}\end{cases}$ (10)
    where $D$ is a constant corresponding to the brightness of a particle and $d_{\text{horiz},n}$ is the pre-determined horizontal distance from the camera where $d_{\text{horiz},-1}=0$. This helps to reduce the computational cost for the following Gaussian filtering.
    (a) Enlarged potion
    (b) 3D plot
    (c) Enlarged potion
    (d) 3D plot
    Figure 4: Marine snow examples in real underwater images. Left: Type H. Right: Type V.
    Figure 5: Illustrative example of marine snow particles. Particles are synthesized at the blue dot locations.
  2. ii.
    Gaussian filtering to particles: Let $g(x,y;\sigma_{n})$ be a Gaussian filter centered at $(x,y)$ with the parameter $\sigma_{n}$. We apply $g(x,y;\sigma_{n})$ to $E_{n}$ for yielding $S_{n}$ for the prototype of the marine snow artifacts.
    $S_{n}(x,y)=g(x,y;\sigma_{n})*E_{n}(x,y).$ (11)
  3. iii.
    Thresholding: To avoid overshooting of the luminance of marine snow artifacts, thresholding is performed as
    $H_{n}(x,y)=\begin{cases}a_{n}S_{n}(x,y)&S_{n}(x,y)<\text{threshold}\\ \text{constant}&S_{n}(x,y)\geq\text{threshold},\end{cases}$ (12)
    where $a_{n}$ is a pre-determined constant to strengthen or weaken the luminance of the artifacts.
    Finally, an image only containing marine snow artifacts is obtained by $H=\sum_{n}H_{n}$. The synthesized type H marine snow artifact is shown in Figs. 6(a) and (b), which are similar to the real version in Figs. 4(a) and (b).
We model the type V artifacts by applying the Laplacian filter to the type H artifacts. For type V, steps i) and ii) are the same as type H. Step iii) is different as:
  • iii’)
    Laplacian filtering and thresholding:
    $V_{n}(x,y)=\begin{cases}L(\sigma_{n})*{a^{\prime}_{n}S^{\prime}_{n}(x,y)}&V<\text{threshold}\\ \text{constant}&V\geq\text{threshold},\end{cases}$ (13)
    where $L(\sigma_{n})$ is a Laplacian filter for edge enhancement with the parameter $\sigma_{n}$ controlling sharpness and $a^{\prime}_{n}$ is similar to $a_{n}$ in (12).
    Finally, an image only containing marine snow artifacts is obtained by $V=\sum_{n}V_{n}$. The synthesized type V marine snow artifact is shown in Figs. 6(c) and (d). Like the type H, it is also similar to the artifacts in the real-world image in Figs. 4(c) and (d).
III-D Synthesizing Marine Snow Artifacts
By using the models introduced in Section III-C, we synthesize the marine snow artifacts with underwater images or synthesized underwater images.
Let $U_{c}\in\mathbb{R}^{H\times W\times 3}$ be an underwater image obtained by the model introduced in Section III-A. To synthesize marine snow artifacts, we simply overlay $H(x,y)$ and $V(x,y)$ on $U_{c}(x,y)$ as:
$I_{c}(x,y)=U_{c}(x,y)+H(x,y)+V(x,y),$ (14)
where $H(x,y)$ and $V(x,y)$ are images with the type H and V marine snow artifacts shown in (12) and (13), respectively.
(a) Enlarged portion
(b) 3D plot
(c) Enlarged portion
(d) 3D plot
Figure 6: Synthesized marine snow artifact examples. Left: Type H. Right: Type V.
IV PHISWID Specifications
Our developed dataset, PHISWID, contains $2264$ image pairs, all having a pixel resolution of $384\times 384$. All original atmospheric RGB-D images used are collected from NYD-RGB dataset [32] and an outdoor image dataset [33]. An image pair contains one original atmospheric image and one synthesized underwater image degraded by color shift and marine snow artifacts. We randomly set water type for each image. Image examples in PHISWID are shown in Fig. 7e.
Each synthesized image contains marine snow artifacts. The number of the type H marine snow particles is randomly generated with a Gaussian distribution of mean 40 and variance 5, and that of the type V marine snow particles conforms to a Gaussian distribution of mean 30 and variance 5.
Most parameters are determined according to our preliminary observations of real underwater images: They are summarized in Appendix B. Further examples of PHISWID images are provided in Appendix C.
Figure 9: Example images in PHISWID.
Figure 7: *
Original images.
Figure 8: *
Synthesized images.
V Benchmarking Results on Synthesized Images
In this section, we compare three datasets (PHISWID, UIEB, and LSUI) using the same module to demonstrate the contribution of our dataset to underwater image enhancement. In addition, we compare other modules using different datasets.
V-A Setup
TABLE IV: Comparison methods.
Abbreviation Baseline module Training dataset
U-Net(P) U-Net PHISWID
U-Net(P’) U-Net PHISWID w/o marine snow
Deep WaveNet(P) Deep WaveNet PHISWID
U-Net(U) U-Net UIEB
Deep WaveNet(U) [43] Deep WaveNet UIEB
WaterNet [11] WaterNet UIEB
U-Net(L) U-Net LSUI
U-shape [37] Transformer LSUI
HLRP [28]
Model-based
enhancement
We divided PHISWID to 2184 training image pairs and 80 test image pairs. Underwater image enhancement performance is compared with methods summarized in Table IV. The parameters of the alternative methods are basically followed to the original researches. We use WaterNet [11], Deep WaveNet [43], and HLRP [28], which are widely used for underwater image enhancement and restoration. We also trained Deep WaveNet with the proposed PHISWID to compare dataset qualities.
Furthermore, we train basic U-Nets with PHISWID, UIEB [11], and LSUI [37]. The structure of the U-Net is shown in Appendix D. For a comparison purpose, we also train the same U-Net with the dataset without marine snow artifacts, i.e., the dataset only applied our color shift model.
V-B Results
The experimental results are shown in Fig. 10 and the objective performances are summarized in Table V. Both PSNRs and SSIMs of the neural networks trained with PHISWID outperform those with the existing datasets. Fig. 10 also visualizes such differences: The methods trained with PHISWID restore images like their atmospheric counterparts. On the other hand, marine snow artifacts remain and blueish images are still generated by methods trained with the alternative datasets. U-Net(P’) can remove bluish color shift, however, it cannot remove marine snow artifacts.
TABLE V: Average PSNRs and SSIMs over the test dataset of PHISWID.
Method PSNR SSIM
U-Net(P) 23.97 0.714
U-Net(P’) 21.68 0.654
Deep WaveNet(P) 28.09 0.818
U-Net(U) 20.05 0.162
Deep WaveNet(U) 21.24 0.166
Water Net 20.89 0.439
U-Net(L) 21.33 0.319
U-shape 24.49 0.413
HLRP 15.59 0.118
Synthesized image 19.51 -0.010
(a) Test images
in PHISWID
(b) U-Net(P)
(c) U-Net(P’)
(d) Deep
WaveNet(P)
(e) U-Net(U)
(f) Deep
WaveNet(U) [43]
(g) WaterNet [11]
(h) U-Net(L)
(i) U-shape [37]
(j) HLRP [28]
Figure 10: PHISWID results.
VI Real-World Examples
In this section, we present image enhancement results by methods trained by PHISWID for real-world underwater images. We also discuss limitations.
VI-A Real-world Underwater Image Enhancement
Fig. 11 shows examples of underwater image enhancement for real-world images. We compare the methods used in Section V-A. As visual evaluation, methods trained with our proposed dataset remove color shift sufficiently. Especially, a significant difference between PHISWID and the other datasets can be observed in background regions. The PHISWID-based methods can remove the bluish color shift in the background while the others fails.
According to the removal of marine snow artifacts, the PHISWID-based methods succeeded remove them. U-Net(P’) and the methods trained with the existing datasets without marine snow artifacts naturally fail to remove the artifacts. This suggests that PHISWID is beneficial for real-world underwater image enhancement.
(a) Original
underwater images
(b) U-Net(P)
(c) U-Net(P’)
(d) Deep
WaveNet(P)
(e) U-Net(U)
(f) Deep
WaveNet(U)
(g) Water Net
(h) U-Net(L)
(i) U-shape
(j) HLRP
Figure 11: Underwater image enhancement results for real-world images.
VI-B Limitations
Although our PHISWID are carefully designed based on physical measurement processes, we still assume many parameters and settings: This is one of the limitations.
As observed in Fig. 11, thick marine snow artifacts still remain in the restored images. The reason for this could be two-fold: First, the architectures of the neural networks are not specifically designed for underwater image enhancement with marine snow artifacts. Second, our proposed model classifies marine snow artifacts into two representative types. Our modeling based on this classification could be incomplete since particles causing marine snow have various structures [58].
With regard to PHISWID, if the light scattering is very different from the training images, color shift, particularly in the distant backgrounds, could not be well restored as shown in Fig. 12. Its possible reason is that the RGB-D images used in PHISWID are recorded in indoor situations which results in limited horizontal distances (i.e., depth) in scenes. Therefore it may not generate color shifts of distant backgrounds.
Our future work includes to update PHISWID for covering wider degradation patterns like hazes.
Figure 12: Limitations. Top: Original images. Bottom: Restoration results by U-Net trained with PHISWID.
VII Conclusions
In this paper, we propose PHISWID for underwater image enhancement. The proposed dataset has a wide variety of underwater image degradation including overlooked effects of marine snow. We physically modeled color shift and marine snow artifacts, and synthesized them in atmospheric RGB-D images. The first benchmarking is performed with PHISWID. The benchmark suggested that methods aided our dataset can improve underwater image enhancement objectively and subjectively.
Appendix A Detailed Derivation of The Jaffe-Macglamery Model
We describe detailed derivations of the Jaffe-MacGlamery model introduced in Section III-C1.
The following is the original equation in the main paper:
$E_{\mathrm{d}}(x,y)=E_{\mathrm{I}}\left(x,y,0\right)e^{-\beta_{c}R}M\left(x,y\right)\cdot\frac{\cos^{4}\theta T_{l}}{4f_{n}}\left[\frac{R-F_{l}}{R}\right]^{2}.$ (15)
By substituting angles into (15), the equation is rewritten as:
$E_{\mathrm{d}}(x,y)=E_{\mathrm{I}}\left(x,y,0\right)e^{-\beta_{c}R}M\left(x,y\right)\cdot\frac{T_{l}}{4f_{n}}\left[\frac{R-F_{l}}{R}\right]^{2},$ (16)
where $E_{\mathrm{I}}\left(x,y,0\right)$ is defined as [36]:
$E_{\mathrm{I}}\left(x,y,0\right)=A^{\prime}\cos\gamma\frac{e^{-\beta_{c}R}}{R^{2}}*p+A^{\prime}\cos\gamma\frac{e^{-\beta_{c}R}}{R^{2}},$ (17)
where $A^{\prime}$ is a constant. Substituting (17) into (16), we obtain:
$E_{\mathrm{d}}(x,y)=$ $(A^{\prime}\frac{e^{-\beta_{c}R}}{R^{2}}*p+A^{\prime}\frac{e^{-\beta_{c}R}}{R^{2}}))e^{-\beta_{c}R}$ (18)
×$\times M\left(x,y\right)\frac{\cos^{4}\theta T_{l}}{4f_{n}}\left[\frac{R-F_{l}}{R}\right]^{2},$
where $M\left(x,y\right)$, $T_{l}$, and $f_{n}$ are constant. By merging all the above constant values, we finally obtain the following:
$E_{\mathrm{d}}(x,y)=(A\frac{e^{-2\beta_{c}R}}{R^{2}}*p+A\frac{e^{-2\beta_{c}R}}{R^{2}})\left[\frac{R-F_{l}}{R}\right]^{2}.$ (19)
Appendix B Parameters for Marine Snow Models
Table VI summarizes the parameters used for modeling marine snow artifacts.
Appendix C Additional Example Images from PHISWID
Fig. 13j show additional examples from PHISWID.
Figure 14: *
Figure 15: Example images in PHISWID.
Figure 13: *
Original images.
Synthesized images.
Appendix D U-Net Structure
The structure of U-Net [38] used for our underwater image enhancement experiments in Sections V and VI are illustrated in Fig. 16. This is basically the same as the original U-Net but we slightly changed the number of convolution layers and that of channels in each layer. The hyperparameters are summarized in Table VII.
Figure 16: U-Net structure.
TABLE VI: Parameter settings for marine snow synthesis.
Type Parameter Setting Corresponding equation
or subsection
H {$d_{-1,\text{horiz}}$, $d_{0,\text{horiz}}$, $d_{1,\text{horiz}}$, $d_{2,\text{horiz}}$} {00, $64$, $128$, $192$} (5)
H $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$, $\sigma_{4}$} {7, 5, 3, 3} Sec. 3.2
H {$a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$} {80, 100,150,200} (6)
H threshold 80 (6)
V {$d_{-1,\text{horiz}}$, $d_{0,\text{horiz}}$, $d_{1,\text{horiz}}$, $d_{2,\text{horiz}}$} {00, $64$, $128$, $192$} (5)
V { $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$, $\sigma_{4}$7, 5, 4, 4} Sec. 3.2
V $\sigma$ 0.2 (7)
V {$a^{\prime}_{1}$, $a^{\prime}_{2}$, $a^{\prime}_{3}$, $a^{\prime}_{4}$} {70, 80, 120, 150} (7)
V threshold 28 (7)
TABLE VII: Parameter settings for U-Net.
Parameter Setting
Learning rate 0.001
Number of epochs 150
Initialized bias values 0